# How Does Perfect Fitting Affect Representation Learning? On the Training Dynamics of Representations in Deep Neural Networks

### v1

@@ 4.3 Analysis of the Linear Classifier Probes Results
In Figs. 4, 5, 18 we show results of the decision regions similarity (DRS) of the linear classifier probes, each diagram corresponds to a different layer in the DNN. For a start, note that the DRS evaluations in Figs. 4, 5 support insights that we got from the CKA in Figs. 1, 2. Specifically, note in Figs. 4a, 5a that towards and in the perfect fitting regime of training, the first layer becomes more similar to random representations as it had at the initialization. Yet, there are two main differences here compared to the CKA results. First, the DRS values have a wider range than in the CKA for the first layer, suggesting that indeed there is a significant training dynamic where the first layer representations become more similar to random representations. Second, the DRS shows that the similarity to random representations towards and in the perfect fitting regime is temporary and eventually at the end of training the first layer representation is different than for random representations. The ViT results in Fig. 18 show a different behavior than ResNet, but this was also the case for the CKA evaluations.
Moreover, comparing the DRS evaluations in Figs. 4d, 4e to Figs. 5d, 5e supports our CKA-based insights that when an epoch-wise double descent occurs, the representations at the deeper layers in the DNN are more actively evolving (i.e., the DRS values at the deeper layers are lower when epoch-wise double descent occurs).
Figure 6: Test error of the linear classifier probes during a ResNet-18 training on CIFAR-10 (a) without label noise, (b) with 20% label noise. Each curve color refer to another layer in the DNN.
Importantly, our linear classifier probes here show that an epoch-wise double descent is reflected only by the linear classifier probes at the deeper layers. See the test error curves of the linear classifier probes as function of the training epoch of the original DNN (Fig. 6). Specifically, in the experiment with label noise (and epoch-wise double descent), only the probe of the fourth ResNet block shows a significant epoch-wise double descent (Fig. 6b). This suggests that the epoch-wise double descent phenomenon crucially depends on the deeper layers and their fine grained features.
@@ 5 Conclusion
In this paper we have elucidated the training dynamics of representations in deep learning. We have used the CKA similarity measure and proposed a new similarity measure based on decision regions of linear classifier probes. Our extensive experimental evaluations provide a highly detailed description of the representation evolution during training. By this, we provide several new insights on the representation learning dynamics, with a particular focus on overparameterized learning that continues much after perfect fitting starts. We believe that the new insights and empirical analysis tools of this paper can pave the way for additional works on related research questions.
@@
Appendix A Limitations
In this paper we study the representational similarity during training of DNNs. We use the CKA similarity measure and propose a new similarity measure based on the decision regions of linear classifier probes. Measuring representational similarity is known as an intricate task, and indeed any similarity measure has its own strengths and weaknesses. Accordingly, while our results provide new insights, most of which are confirmed by the two similarity measures we use, other similarity measures may be proposed in the future and lead to additional insights and discoveries.
Moreover, we present in this paper a highly extensive, computationally-intensive set of experiments on various training settings. Nevertheless, we focus on the overparameterized learning settings where the trained model is able to perfectly fit its training data, and the training continues much after the point where perfect fitting starts; our focus on these learning settings conforms with the research questions that we address in this paper and related to the effect of overparameterization and perfect fitting on the representation learning process. While we provide an exemplary experiment for training with explicit regularization (Adam with weight decay), future work can consider in more detail regularized training settings and search for new insights for them.
Appendix B Broader Impact
Appendix C Experiments Compute Resources
For all of our experiments, we utilized an internal cluster equipped with Nvidia RTX 3090 and RTX 4090 GPUs, each with 24GB of memory. Below is a summary of the compute resources and run times for each experiment:
  • Model training:
    • ResNet-18 and ViT-B/16 models trained on the CIFAR-10 dataset: Approximately 15 hours.
    • ResNet-18 and ViT-B/16 models trained on the Tiny ImageNet dataset: Approximately 48 hours.
  • CKA similarity heatmap diagrams:
    • Generation of CKA similarity heatmap diagrams: Approximately 15 hours per model, using 6 GPUs.
  • Decision region similarity (DRS) of linear classifier probes:
    • Generation of DRS heatmaps: Approximately 3 days per model, using 12 GPUs.
In terms of the number of experiments:
  • Training base models:
    • We trained approximately 10 models of ResNet-18 on CIFAR-10:
        4 different widths (32, 42, 52, 64).
        3 different label noise levels (with width parameter k=64).
        2 models were trained with SGD, with and without label noise (all the other models were trained with the Adam optimizer).
        1 model was trained with weight decay regularization.
    • We trained 2 ResNet-18 models on the Tiny ImageNet dataset, with and without label noise.
    • We trained 1 ResNet-50 models on the CIFAR-10 dataset, with 20% label noise.
    • We trained 2 models of ViT-B/16 on the CIFAR-10 dataset, with and without label noise.
  • CKA similarity heatmap diagrams:
    • For each model we trained, we also produced the CKA heatmap diagram, resulting in a total of approximately 14 CKA heatmap experiments.
  • Decision region similarity (DRS) of linear classifier probes:
    • We produced DRS heatmaps for ResNet-18 with 20% label noise, without label noise, and for ViT-B/16 without label noise, resulting in a total of 3 DRS heatmap experiments.
In terms of storage requirements:
  • For storing ResNet-18 models, we stored a checkpoint of the model after every epoch of the 4000 epochs, requiring approximately 170GB of storage per experiment (for all 4000 epochs).
  • For storing ViT-B/16 models, we stored a checkpoint of the model following this grid: all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999, resulting in a total of 350GB per experiment.
  • In total, we used around 4.5 TB of storage for all experiments described in this paper.
We did not use significant compute resources for any other tasks beyond the specified above.
Appendix D Assets and Software Packages
@@
  • dbViz: For decision region similarity (DRS) experiments, we used methods introduced in [12] with code available at https://github.com/somepago/dbViz, licensed under the Apache 2.0 License.
Appendix E Additional Related Work
In [9], the authors propose to analyze the first layer in convolutional DNNs using the energy spectrum of the learned filters in the PCA basis of the input image patches. Their results show a typical form of the learned energy spectrum which is consistent over different convolutional architectures and datasets but very different compared to its random initialization at the beginning of training; we disprove the generality of this claim in this paper, by using the CKA similarity metric and showing that the representations of the first layer of a ResNet model can be very similar to the representations at the random initialization. This difference is probably due to our training setting that allows long training in the perfect fitting regime and avoids early stopping (whereas their training is much shorter; for example, their Fig. 18c in [9] shows a training process of 200 epochs that stops approximately when perfect fitting starts).
In [10], the authors use synthetic datasets (with controllable feature relevance) to study which features a trained model uses and how the choice of the features relates to their relevance to the task and to their linear-decodability from representations in an untrained model. Moreover, they show that trained models have relatively higher representational similarity (quantified by the neuroscientific method from [13]) to untrained models than to models trained on another task. Importantly, the representational similarity between trained and untrained models which is reported in [10] is moderate, and even quite low, in contrast to our results in this paper that shows much higher similarity levels according to other similarity metrics.
Appendix F Additional Details on the Experiment Settings
F.1 Models
ResNet-18: We used the (Preactivation) ResNet-18 architecture as introduced in [14]. The ResNet-18 structure includes an initial convolutional layer followed by four ResNet blocks, each consisting of two BatchNorm-ReLU-Convolution layers. The layer widths for these four ResNet blocks are defined as [k, 2k, 4k, 8k] where $k$ is a width parameter (positive integer), and $k=64$ corresponds to the standard ResNet-18 layout. Similarly to the deep double descent research in [1], we consider several values of $k$ to examine the effect of width and parameterization level on the ResNet-18 model. The ResNet-18 implementation was as in [1, 12] that originate in https://github.com/kuangliu/pytorch-cifar.
ResNet-50: For the experiment on ResNet-50 in Fig. 12 we used a standard ResNet-50 implementation that originates in https://github.com/kuangliu/pytorch-cifar and follows [14]. No changes were applied on the model, including not on its standard width.
ViT-B/16: In our implementation, we utilized the Vision Transformer (ViT) architecture, specifically the ViT-B/16 model, as described in [15]. The ViT-B/16 structure begins with an initial patch embedding layer, followed by a sequence of 12 transformer blocks. We used the ViT-B/16 implementation of PyTorch.
F.2 CKA Experimental Details
@@
The CKA evaluations considered the following epoch grids:
  • For ResNet-18 and ResNet-50 training on CIFAR-10, the epoch grid is always full with all the epochs from 0 to 3999.
  • For ResNet-18 training on Tiny ImageNet, the epoch grid includes all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999.
  • For ViT-B/16, the epoch grid includes all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999.
F.3 Decision Regions Similarity (DRS) Experimental Details
@@
    We randomly chose 500 triplets of training examples (these triplets were constant for all DRS experiments). Then, for each triplet, we computed the plane that they span (as mathematically explained in [12] and based on their code). The $j^{\text{th}}$ plane spanned by those 3 samples, denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$, is a 2D plane in the input space. We then discretized each plane $\mathcal{Q}_{j}$ into a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\text{discrete}}|=2500$ points.
@@
Appendix G Summary of Experiment Settings
@@
  • CKA experiments: Please refer to Table 1 for a summary of our CKA experiments, describing the datasets, architectures, optimizers, width parameters $k$, label noise percentage, and the corresponding figures of the results.
@@
  • DRS experiments: As detailed in Section C, generating the Decision Region Similarity (DRS) heatmaps is highly computationally demanding, requiring significant time and compute resources. Therefore, we limited the DRS evaluations to the following configurations:
    • ResNet-18 trained on CIFAR-10 without label noise (Fig. 4).
    • ResNet-18 trained on CIFAR-10 with 20% label noise (Fig. 5).
    • ViT-B/16 trained on CIFAR-10 without label noise (Fig. 18).
Dataset Architecture Optimizer Width P. ($k$) % Label Noise Fig. #
CIFAR-10 ResNet-18 Adam 64 0 1
10 7
20 2
52 20 10
42 20 9
32 20 8
SGD+Mom. 64 20 14
SGD+Mom. 64 0 13
SGD (0.01) 64 20 15
Adam (wd) 64 0 16
ResNet-50 Adam N.A 20 12
ViT-B/16 Adam N.A 0 3
ViT-B/16 Adam N.A 20 17
Tiny ImageNet ResNet-18 Adam 64 0 11
20 20
Tiny ImageNet $1/2$ ResNet-18 Adam 64 20 21
@@
Table 1: Summary of the CKA experiments conducted. Note that here Adam (wd) stands for training with Adam optimizer with weight decay regularization of 0.001, and Tiny ImageNet $1/2$ stands for the experminet where we used only half of the train samples of Tiny ImageNet. In addition, SGD (0.01) refers to Fig. 15, where we used SGD optimizer with constant learning rate of 0.01 and no momentum.
Appendix H Additional Experimental Results
Figure 7: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 10% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.
Figure 8: CKA evaluations for ResNet-18 of width parameter $k=32$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 9: CKA evaluations for ResNet-18 of width parameter $k=42$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 10: CKA evaluations for ResNet-18 of width parameter $k=52$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 11: CKA evaluations for ResNet-18 (of standard width), trained on Tiny ImageNet (image size $32\times 32\times 3$) without label noise.
@@ Appendix H Additional Experimental Results
Figure 12: CKA evaluations for ResNet-50, trained on CIFAR-10 with 20% label noise.
Figure 13: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Here, the optimizer is SGD with constant learning rate 0.0001 and momentum of 0.9
Figure 14: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is SGD with constant learning rate 0.0001 and momentum of 0.9
Figure 15: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is SGD with constant learning rate 0.1 without momentum
Figure 16: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is Adam with weight decay of 0.001.
Figure 17: CKA evaluations for ViT-B/16, trained on CIFAR-10 with 20% label noise.
Figure 18: Decision regions similarity (DRS) of linear classifier probes. Evaluations for ViT-B/16, trained on CIFAR-10 without label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The DNN error curves are provided in Fig. 3g.
H.1 CKA Similarity of the First Layer Representations to Random Representations
As mentioned in Section 3.3 of the main text, the similarity of the first layer representations to the random representations (as appear, e.g., in Figs. 2a, 1a) is not only to the specific random initialization of the training process, but also to other random representations unrelated to the specific training process. This can be observed in Fig. 19 where we evaluate the CKA similarity of the first layer representations during training with respect to a random representation that was not used as the initialization of the specific training process; yet, it should be noted that the CKA similarity to the specific initialization is 1%-3% higher than the CKA similarity to another initialization.
Figure 19: First row from the CKA diagram in Fig. 2a for the first convolutional layer (blue line), together with the CKA similarity with respect to another random initialization (orange line). The model is ResNet-18 trained on CIFAR-10 with 20% label noise.
H.2 Additional CKA Evaluations
In Figs. 7-17 we provide CKA evaluations for various learning settings, in addition to the results in the main paper. See Table 1 for the summary of training settings.
H.3 Additional DRS Evaluations
In Fig. 18 we provide DRS evaluation for the ViT-B/16 model. This is in addition to the DRS evaluations on ResNet-18 in the main paper.
@@ H.3 Additional DRS Evaluations
Figure 20: CKA evaluations for ResNet-18 (of standard width), trained on Tiny ImageNet with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.

### v2

@@ 4.2.2 Analysis of the DRS Results
In Figs. 3, D.29 -D.34 we show DRS diagrams, each corresponds to a different layer in the DNN. For a start, note that the DRS evaluations in Figs. 3, D.29 support insights that we got from the CKA in Figs. 1, D.1. Specifically, recall Section 3.3.3 and note in Figs. 3a, D.29a that towards and in the perfect fitting phase of ResNet training, the first layer becomes more similar to random representations as at the initialization. Yet, there are two main differences here compared to the CKA results. First, the DRS values have a wider range than in the CKA for the first layer, suggesting that indeed there is a significant training dynamics of task-specific features in the first layer. Second, the DRS shows that the similarity to random representations towards and in the perfect fitting phase is temporary and later when the training ends the first layer representations may differ from random representations.
In Figs. 4, 5, F.1, F.2, F.3, F.4 we visualize the evolution of decision regions during training at different layer probes, as well as the DNN prediction output. Each of the visualized planes shows the geometrical evolution of the decision regions, demonstrating when during training the decisions change more, less or even stabilize (at least approximately). These visualizations also show geometrical properties of the decision regions, e.g., decision regions size and fragmentation level (i.e., number of fragments per visualized plane segment). In Figs. E.2, E.2 we complement the visualizations by providing quantitative evaluations of the average number of fragments per plane at various layers and DNN output during training (see Appendix E for more details).
In Figs. 4, 5, F.1, F.2, F.3, F.4 we observe that the decision regions evolve significantly different depending on the optimizer and DNN architecture – reflecting the different implicit regularization that different optimizers and architectures may have. In Figs. 4, F.1, F.2, for Adam, we observe the first layer behavior where late in training the decision regions approximately revert to their random initialization form; in contrast, in SGD, the first layer evolution includes two main phases where late in training the decision regions stabilize and differ from their initialization.
The quantitative fragmentation evaluations (Figs. E.2, E.2) and the visualizations (Figs. 5, F.3) show that the decision regions of the DNN output and the deeper-layer probes may become more fragmented during the training phase of memorizing atypical examples (without artificially added label noise), and then the fragmentation level stabilizes in the perfect fitting training phase (possibly with some reduction in fragmentation before the stabilization). This behavior is observed more clearly for ViT, also due to its synchronized training dynamics across its layers — emphasizing a significant aspect of implicit regularization on representation learning dynamics in ViT compared to ResNet.
@@ 5 Conclusion
In this paper we have elucidated the training dynamics of representations in deep learning. We have used the CKA similarity measure and proposed a new similarity measure based on decision regions of linear classifier probes. Our extensive experimental evaluations provide a highly detailed description of the representation evolution during training. By this, we provide new insights on representation learning dynamics and how they depend on the architecture (ResNet vs. ViT) and optimizer (Adam vs. SGD). We believe that the new insights and empirical analysis tools of this paper can pave the way for more studies on related research questions.
Broader Impact
@@
Appendix A Related Work: Additional Discussion
The effect of DNN width on similarity of representations was examined by Morcos et al. (2018) using the CCA metric, and by Kornblith et al. (2019); Nguyen et al. (2021) using the CKA metric. In contrast to these works, our focus is on elucidating the representational similarity during the training process. In the following, we discuss the existing literature which is more related to our work.
A.1 Representational Similarity Throughout the Training Process
Raghu et al. (2017) proposed the SVCCA metric for representational similarity and used it for several analyses, including showing how the representational similarity between the model during training and the model at the end of training evolves during training of a convolutional model and a ResNet model. In addition, Liu et al. (2021) showed the evolution of the SVCCA metric during 4 epochs of a BERT model training. In this paper, we examine the evolution of representational similarity much more extensively than previous works, we consider various new settings (such as for a ViT architecture) and, most importantly, we use it for the first time as a tool for analyzing training that continues much after perfect fitting.
Gotmare et al. (2019) examined the loss landscape (specifically, connectivity of local minima) and representational similarity (using the CCA metric (Raghu et al., 2017)) for various training heuristics from the common practice, such as learning rate restarts, warmup, and distillation. Their results include evaluation of CCA representational similarity between a VGG model at training iteration 0 and the specific training iteration at the end of their warmup phase (if applied); their evaluation reflects high similarity for the first convolutional layer of the model. They also evaluate the CCA similarity between the model at the end of training and at the intermediate training iteration after the warmup. Importantly, the work by Gotmare et al. (2019) does not show or analyze the detailed evolution of the representational similarity during the entire training process. In contrast, in our paper here we examine the representational similarity in much more detail, enabling to understand the representational similarity of various pairs of epochs, throughout a long training process with a small constant learning rate that does not affect the natural training dynamics. Moreover, we focus here on understanding how the representational similarity evolves in overparameterized learning where perfect fitting does not prevent good generalization.
A.2 Representational Similarity between Trained and Untrained Models
Kornblith et al. (2019) examined the CKA similarity between trained and untrained convolutional DNNs. They showed that the representations in the first layers of the DNN are similar to those of the corresponding layers in the untrained DNN. Their CKA values for the first layer of a 10-layer convolutional model on CIFAR-10 was slightly above 0.8. In this paper (e.g., Figs. 1a, D.1a, D.2a), we show that the first layer can have CKA similarity (nearly) 1 in the perfect fitting regime of training.
Chowers & Weiss (2023) propose to analyze the first layer in convolutional DNNs using the energy spectrum of the learned filters in the PCA basis of the input image patches. Their results show a typical form of the learned energy spectrum which is consistent over different convolutional architectures and datasets but very different compared to its random initialization at the beginning of training; we disprove the generality of this claim in this paper, by using the CKA similarity metric and showing that the representations of the first layer of a ResNet model can be very similar to the representations at the random initialization. This difference is probably due to our training setting that allows long training in the perfect fitting regime and avoids early stopping (whereas their training is much shorter; for example, their Fig. 18c in (Chowers & Weiss, 2023) shows a training process of 200 epochs that stops approximately when perfect fitting starts).
Hermann & Lampinen (2020) use synthetic datasets (with controllable feature relevance) to study which features a trained model uses and how the choice of the features relates to their relevance to the task and to their linear-decodability from representations in an untrained model. Moreover, they show that trained models have relatively higher representational similarity (quantified by the neuroscientific method of Kriegeskorte et al. (2008)) to untrained models than to models trained on another task. Importantly, the representational similarity between trained and untrained models which is reported by Hermann & Lampinen (2020) is moderate, and even quite low, in contrast to our results in this paper that shows much higher similarity levels according to other similarity metrics.
Appendix B Additional Details on the Experiment Settings
B.1 Software Packages
@@
  • dbViz: For decision region similarity (DRS) experiments, we used methods introduced in (Somepalli et al., 2022) with code available at https://github.com/somepago/dbViz, licensed under the Apache 2.0 License.
B.2 Models
ResNet-18: We used the (Preactivation) ResNet-18 architecture as introduced in (He et al., 2016). The ResNet-18 structure includes an initial convolutional layer followed by four ResNet blocks, each consisting of two BatchNorm-ReLU-Convolution layers. The layer widths for these four ResNet blocks are defined as [k, 2k, 4k, 8k] where $k$ is a width parameter (positive integer), and $k=64$ corresponds to the standard ResNet-18 layout. Similarly to the deep double descent research in (Nakkiran et al., 2021), we consider several values of $k$ to examine the effect of width and parameterization level on the ResNet-18 model. The ResNet-18 implementation was as in (Nakkiran et al., 2021; Somepalli et al., 2022) that originate in https://github.com/kuangliu/pytorch-cifar.
ResNet-50: For the experiment on ResNet-50 in Fig. D.7 we used a standard ResNet-50 implementation that originates in https://github.com/kuangliu/pytorch-cifar and follows (He et al., 2016). No changes were applied on the model, including not on its standard width.
ViT-B/16: In our implementation, we utilized the Vision Transformer (ViT) architecture, specifically the ViT-B/16 model, as described in (Dosovitskiy et al., 2020). The ViT-B/16 structure begins with an initial patch embedding layer, followed by a sequence of 12 transformer blocks. We used the ViT-B/16 implementation of PyTorch.
B.3 CKA Experimental Details
@@
The CKA evaluations considered the following epoch grid - all epochs from 0 to 300, every third epoch between 300 to 900, and every fifth epoch from 900 to 3999.
B.4 Atypical Inputs Setting
In selected experiments on datasets without artificially added label noise, we evaluate the DNN classification erroron atypical examples from its train dataset. The atypical examples are found as follows:
  • CIFAR-10: We use the C-score defined by Jiang et al. (2021) and set the threshold to $0.5$. All images with $\mathrm{C\text{-}score}>0.5$ are considered atypical.
  • SVHN: We use the loss-based C-score introduced by Wu et al. (2020), adjusting its loss threshold to match the proportion of atypical inputs found for CIFAR-10.
B.5 Decision Regions Similarity (DRS) Experimental Details
@@
    We randomly chose 500 triplets of training examples (these triplets were constant for all DRS experiments). Then, for each triplet, we computed the plane that they span (as mathematically explained in (Somepalli et al., 2022) and based on their code). The $j^{\text{th}}$ plane spanned by those 3 samples, denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$, is a 2D plane in the input space. We then discretized each plane $\mathcal{Q}_{j}$ into a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\text{discrete}}|=2500$ points.
@@
Appendix C Summary of Experiment Settings
@@
  • CKA experiments: Please refer to Table C.1 for a summary of our CKA experiments, describing the datasets, architectures, optimizers, width parameters $k$, label noise percentage, and the corresponding figures of the results.
@@
  • DRS experiments: Generating the Decision Region Similarity (DRS) heatmaps is highly computationally demanding, requiring significant time and compute resources. Therefore, we limited the DRS evaluations to only a subset of the of the experiment settings. Please refer to Table C.1 for the relevant figures.
Dataset Architecture Optimizer Width P. ($k$) % Label Noise CKA Fig. # DRS Fig. #
CIFAR-10 ResNet-18 Adam 64 0 1 3
10 D.2
20 D.1 D.29
52 20 D.5
42 20 D.4
32 20 D.3
SGD (0.001) 64 0 D.15
SGD+Mom. 64 20 D.9
SGD+Mom. 64 0 D.8
SGD (0.01) 64 20 D.10
Adam (wd) 64 0 D.11
ResNet-50 Adam N.A 20 D.7
ViT-B/16 Adam N.A 0 2 D.30
ViT-B/16 SGD (0.001) N.A 0 D.22 D.31
ViT-B/16 Adam N.A 20 D.12
Tiny ImageNet ResNet-18 Adam 64 0 D.6
20 D.14
Tiny ImageNet $1/2$ ResNet-18 Adam 64 20 D.16
SVHN ResNet-18 Adam 64 0 D.18 D.32
SGD (0.001) 64 0 D.19 D.33
ViT-B/16 Adam N.A 0 D.23
SGD (0.001) N.A 0 D.24
@@
Table C.1: Summary of the CKA experiments conducted. Note that here Adam (wd) stands for training with Adam optimizer with weight decay regularization of 0.001, and Tiny ImageNet $1/2$ stands for the experminet where we used only half of the train samples of Tiny ImageNet. In addition, SGD (0.01) refers to Fig. D.10, where we used SGD optimizer with constant learning rate of 0.01 and no momentum.
Appendix D Additional Experimental Results
@@
Figure D.1: CKA evaluations for ResNet-18 trained on CIFAR-10 with 20% label noise.
@@
Figure D.2: CKA evaluations for ResNet-18 trained on CIFAR-10 with 10% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.